TheBerkeleyUNIXNetwork EricSchmidt OverallIntroduction Thissetofpapersdocumentsanetworkconstructedtosatisfytheproject optionfortheMaster’sdegreefromtheEECSDepartment,C.S.Division,U.C. Berkeley. Thefirstpaperisatechnicaldescriptionofthehistoryoftheproject,details ofimplementation,andsummarypoints.Thesecondpaperisanintroductionbeing soldtomanyUNIX◯rusersontheBerkeleycampus.Thethirdgroupofpagesare theUNIXProgrammersManualsectionsforthenetworkcommandsavailable.The lastpaperisamanualforsystemsstaffaboutoperationandsetupofthenetwork. EricSchmidt June1979